# Graph Memory Learning: Imitating Lifelong Remembering and Forgetting of Brain Networks

### v1

@@ Abstract
Graph data in real-world scenarios undergo rapid and frequent changes, making it challenging for existing graph models to effectively handle the continuous influx of new data and accommodate data withdrawal requests. The approach to frequently retraining graph models is resource intensive and impractical. To address this pressing challenge, this paper introduces a new concept of graph memory learning. Its core idea is to enable a graph model to selectively remember new knowledge but forget old knowledge. Building on this approach, the paper presents a novel graph memory learning framework - Brain-inspired Graph Memory Learning (BGML), inspired by brain network dynamics and function-structure coupling strategies. BGML incorporates a multi-granular hierarchical progressive learning mechanism rooted in feature graph grain learning to mitigate potential conflict between memorization and forgetting in graph memory learning. This mechanism allows for a comprehensive and multi-level perception of local details within evolving graphs. In addition, to tackle the issue of unreliable structures in newly added incremental information, the paper introduces an information self-assessment ownership mechanism. This mechanism not only facilitates the propagation of incremental information within the model but also effectively preserves the integrity of past experiences. We design five types of graph memory learning tasks: regular, memory, unlearning, data-incremental, and class-incremental to evaluate BGML. Its excellent performance is confirmed through extensive experiments on multiple real-world node classification datasets.
@@ I Introduction
Nowadays, a multitude of pivotal datasets in practical applications are expressed as graphs [1], encompassing diverse domains such as social networks [2, 3], traffic networks [4, 5], biological networks [6, 7], recommendation systems [8, 9]. Graph learning models attract increasing attention and demonstrate robust data mining and representation learning capacities. However, graph learning faces the challenge of promptly assimilating new information while discarding obsolete knowledge that has been invalidated. This is attributed to the fact that real-world data often exhibits dynamic flow characteristics, wherein a continuous influx and withdrawal of data occur. As depicted in Fig. 1 (A), the flow of data across the graph frequently induces alterations in the overarching topology, thereby exacerbating the difficulties encountered by graph models. Consequently, the objective of this study is to explore and empower graph learning models with the capability to “memory” based on the dynamic graph. This paradigm is named as graph memory learning (GML).
@@
Existing graph lifelong learning (GLL) [10, 11] and graph unlearning (GUL) [12, 13] approaches have undertaken explorations into the memory learning of graph models from limited perspectives. GLL aims to address the challenges of continuous data increments, adapts to new knowledge, and mitigates the issue of forgetting past experiences [14, 15]. GUL emphasizes the right of data to be forgotten, aiming to help the model forget the specified revocation data at the lowest possible cost while minimizing damage to the existing knowledge of the model [16]. However, GML is far from a combination of GLL and GUL, and it entails comprehensive and nuanced contemplation of memory retention and forgetting corresponding to the continuous influx and withdrawal of data, respectively. The inherent conflict between memory retention and forgetting presents a natural barrier that obscures graph learning models from accurately perceiving changes in graph structures, leading to a decrease in their expressive capacity.
@@
In light of the above observations, this paper proposes a Brain-inspired Graph Memory Learning algorithm (BGML) based on brain network dynamics and function-structure coupling. It aims to solve the problem of GML for dynamically evolving graph data. Specifically, this paper proposes a multi-granular feature graph learning scheme with hierarchical progressive learning to clarify changes in local graph knowledge. Key components comprise a multi-granular graph partition (MGP) and a multi-granular hierarchical progressive learning mechanism (MGHPL) based on feature graph grain learning (FGGL). Notably, feature graph grains implicitly encode the graph structure, liberating nodes to operate as independent samples and drawing inspiration from feature graph networks. Furthermore, an information self-assessment ownership mechanism (ISAO) is proposed to tackle the issue of unreliable new knowledge structures. Essentially, this paper explores three questions: (1) What GML approach can empower graph models to selectively remember important information and forget irrelevant information? (2) How to alleviate the conflict between remembering and forgetting in the graph memory model? (3) How to assign a credible graph structure to knowledge during the process of remembering and forgetting? The main contributions are summarized below:
@@
  • Extensive experiments on five tasks targeting different aspects of GML on the node classification task on five real-world datasets demonstrate that BGML outperforms all baselines significantly.
@@ II-A Graph Neural Networks
Graph neural networks (GNNs) have become a prominent method for processing unstructured data [20, 21, 22]. Specifically, Kpif et al. [23] introduced GCNs, utilizing first-order truncated Chebyshev polynomials to approximate graph convolutions. Additionally, Veličković et al. [24] introduced Graph Attention Networks (GATs), incorporating an attention mechanism on the graph, resulting in significant performance improvements. Furthermore, Hamilton et al. [25] were the first to identify the latent nodes within a graph, leading to the proposal of the inductive spatial graph neural network known as GraphSAGE. As research progresses, scholars have observed that the modeling of real-life scenes using graphs typically undergoes dynamic evolution, extending beyond static data [26, 27]. Based on distinct data forms (discrete and continuous), current research can be categorized into two groups [28]: discrete dynamic graph neural networks (DTDG), exemplified by EvolveGCN [29] and DyTed [30], and continuous dynamic graph neural networks (CTDG), exemplified by Ada-DyGNN [31] and APAN [32].
@@ II-B Lifelong Learning and Graph Lifelong Learning
Recently, these strategies have been progressively extended into the domain of graph learning [10]. Regarding architectural design, Wang et al. [14] devised the FGNs method to translate the graph structure into a format compatible with conventional data learning structures. To enhance the encoding of attribute information and the topological structure of target nodes, Zhang et al.’s [15] HPNs extract various levels of knowledge abstraction represented as atoms. Regarding the rehearsal scheme, Zhou et al. [40] store past knowledge in an experience buffer and conduct lifelong learning on graph data through experience replay. In terms of regularization, Kou et al. [41] introduced decoupling-based continuous graph representation learning (DICGRL). tan et al. [42] introduced GPIL with transferable meta-knowledge for initialization learning by introducing a pseudo-incremental approach. Furthermore, this field has given rise to several hybrid methods. ContinualGNN, proposed by Wang et al. [43], combines two approaches to alleviate catastrophic forgetting and preserve the existing learning mechanism.
Existing GLL methods focus on addressing catastrophic forgetting, but they are impractical for limited resources as they aim to retain all information indefinitely. BGML delves into making the model selectively remember important information and forget what is not.
@@ II-C Machine Unlearning and Graph Unlearning
In long-term continuous learning, ensuring model performance and safety requires addressing the critical aspect of enabling the machine to learn to forget [12, 13]. Cao et al. [44] were pioneers in introducing the concept of unlearning. Subsequently, researchers have proposed a series of new methods in two primary directions: exact unlearning and approximate unlearning [45, 46]. SISA (Sharded, Isolated, Sliced, and Aggregated), proposed by Bourtoule et al. [47], stands as one of the most representative works in this regard. The fundamental concept involves randomly dividing the training set into several disjoint shards and training the corresponding shard models independently. Upon receiving the forget request, the model only needs to retrain the respective shard model.
Recently, Chen et al. [16] extended the concepts of SISA to the graph domain, presenting GraphEraser as the inaugural graph unlearning model. Building upon this foundation, Li et al. [48] introduced MEGU to chieve a balance between forgetting performance and framework generalization. Current research on unlearning is still mainly focused on static graphs, and no new issues have been added for linkage and thinking. BGML intends to apply this mechanism to dynamic evolution graphs, representing a bold endeavor.
@@ III Problem Formulation
The main target of GML is the dynamic evolution graph based on discrete snapshots, i.e., $\mathcal{G}=(\mathcal{G}_{0},\mathcal{G}_{1},\mathcal{G}_{2},...,\mathcal{G}_{t})$. The graph evolution process of each snapshot mainly includes two aspects. First, new data entering the graph network can be represented as $\mathcal{G}_{t}^{{}^{\prime}}=\mathcal{G}_{t-1}+\bigtriangleup\mathcal{G}_{t}^{{}^{\prime}}=\mathcal{G}_{t-1}\cup\left\{\left\{v_{\mathrm{new}},e_{\mathrm{new},u}\mid\forall v_{u}\in\mathcal{N}(v_{\mathrm{new}})\right\}\mid v_{\mathrm{new}}\in\mathcal{V}_{\mathrm{add}}\right\}$, where $\mathcal{N}(v_{\mathrm{new}})$ represents the neighbors of $v_{\mathrm{new}}$ and $\mathcal{V}_{\mathrm{add}}$ represents the set of new nodes. Second, when the graph network receives forgetting requests (FR), the clearing operation can be expressed as $\mathcal{G}_{t}^{{}^{\prime\prime}}=\mathcal{G}_{t-1}+\bigtriangleup\mathcal{G}_{t}^{{}^{\prime\prime}}=\mathcal{G}_{t-1}\setminus\left\{\left\{v_{\mathrm{forget}},e_{\mathrm{forget},u}\mid\forall v_{u}\in\mathcal{N}(v_{\mathrm{forget}})\right\}\mid v_{\mathrm{forget}}\in\mathcal{V}_{\mathrm{delete}}\right\}$, where $\mathcal{V}_{\mathrm{delete}}$ is used to represent the set of nodes that need to be forgotten. In general, the change of the graph at time $t$ can be expressed as $\mathcal{G}_{t}=\mathcal{G}_{t-1}+\bigtriangleup\mathcal{G}_{t}$, where $\bigtriangleup\mathcal{G}_{t}=\bigtriangleup\mathcal{G}_{t}^{{}^{\prime}}+\bigtriangleup\mathcal{G}_{t}^{{}^{\prime\prime}}$. In addition, task context changes over time throughout the process of graph memory learning. The task set can be represented by $\mathcal{T}=(\mathrm{T}_{0},\mathrm{T}_{1},\mathrm{T}_{2},...,\mathrm{T}_{t})$. The corresponding time is $t=(t_{0},t_{1},t_{2},...,t_{t})$.
@@ IV Graph Memory Learning Framework
GML essentially focuses on three aspects of learning in dynamic evolution graphs: evolving graph comprehension, learning new information, and forgetting specific information. The modular and hierarchical structure of the brain, as well as the dynamic integration mechanism based on brain network dynamics, inspire solutions to these challenges. Based on these inspirations, this paper introduces a graph memory learning framework (BGML), as shown in Fig. 2.
@@ IV-B MGP: Multi-granular Graph Partition
Brain-Inspiration. The sensory system (e.g., vision, hearing, touch, and smell) provides knowledge to the brain for learning. Various sensory organs receive distinct knowledge stimuli, convert them into neural signals, and transmit them to the brain through diverse neural pathways. Upon reaching the brain, neural signals are allocated to different subdivisions of the cerebral cortex for processing. The cerebral cortex subdivisions (e.g., frontal, parietal, and occipital) govern various perceptual and cognitive functions. Consider the process of distinguishing and remembering fruit types in the brain as an example. The human eye observes the color and shape of the fruit, generating a stimulus signal. Upon transmission to the brain, the occipital lobe area of the cerebral cortex generates a response. Likewise, tactile signals resulting from the aroma and skin characteristics of fruits elicit responses in the parietal.
Additionally, cortical subdivisions of the brain are commonly characterized in neuroscience as exhibiting a multilevel organization. The initial subdivision, responsible for receiving and processing basic sensory information, does not directly facilitate the brain’s decision-making process. Secondary partitions can offer a substantial wealth of detailed information to bolster more sophisticated cognitive functions. For instance, when distinguishing between tangerines and oranges, attention to local details such as the fruit skin at various levels and angles is necessary.
MGP Method. In conclusion, the cerebral cortex’s strategy of multi-region collaboration and hierarchical progression fosters the agent’s profound comprehension of knowledge. This necessitates preprocessing before knowledge input to facilitate problem division and resolution. Inspired by this, the paper devised an MGP algorithm, As shown in the MGP module in Fig. 2. This algorithm aims to partition the static regular graph $\mathcal{G}_{0}$ at time $t_{0}$ into multiple granules, establishing a hierarchy of knowledge details. Subsequent graph evolution details will be simultaneously incorporated into the granule. This lays the groundwork for the subsequent in-depth understanding and hierarchical extraction of precise information in the dynamic evolution graph. Specifically, the MGP algorithm is constructed based on common community detection-based label propagation algorithms (LPA) [49, 50] or $k$-means clustering algorithms (KM) [51]. It is worth mentioning that to mitigate uneven partitioning, the paper employs the balanced versions (BLPA and BEKM) improved by Chen et al. [16]. For $\mathcal{G}_{0}$, the two graph partitions of coarse and fine granularity can be expressed by Eq.1:
@@
where $i=1,2,\dots,k$. $\mathcal{S}_{i}$ represents a first-level graph shard with coarse granularity. $\mathcal{P}_{j}^{i}$, on the other hand, denotes a second-level graph shard within $\mathcal{S}_{i}$, offering finer granularity, where $j=1,2,\dots,l$. $\mathrm{BLPA}(\cdot)$ and $\mathrm{BEKM}(\cdot)$ in Eq.1 are abstract representations of iterative algorithms. Before iteration, they need to be provided with random initialized partition shards $\left\{\mathcal{S}_{1},\mathcal{S}_{2},\dots,\mathcal{S}_{k}\right\}^{0}$ and centroids $\left\{\mathbf{m}_{1},\mathbf{m}_{2},\dots,\mathbf{m}_{k}\right\}^{0}$, respectively. Furthermore, the balance of partitioning is reflected in the fact that, regardless of the partition method employed, the number of nodes within the shards cannot surpass the configurable threshold $\delta$.
BLPA Partition Method. For $\mathrm{BLPA}(\cdot)$, the core iteration steps in which $\forall v\in\mathcal{S}_{i}$ participates can be described as:
@@
BEKM Partition Method. On the other hand, $\mathrm{BEKM}(\cdot)$ does not start directly from the graph structure and needs to obtain the node embeddings $\left\{\mathbf{b}_{1},\mathbf{b}_{2},\dots,\mathbf{b}_{N}\right\}$ through pre-training in advance. It performs clustering by distance between node embedding and centroids. For $\forall v$, the specific iteration steps can be expressed by Eq.3:

### v2

@@ Abstract
Graph data in real-world scenarios undergo rapid and frequent changes, making it challenging for existing graph models to effectively handle the continuous influx of new data and accommodate data withdrawal requests. The approach to frequently retraining graph models is resource intensive and impractical. To address this pressing challenge, this paper introduces a new concept of graph memory learning. Its core idea is to enable a graph model to selectively remember new knowledge but forget old knowledge. Building on this approach, the paper presents a novel graph memory learning framework - Brain-inspired Graph Memory Learning (BGML), inspired by brain network dynamics and function-structure coupling strategies. BGML incorporates a multi-granular hierarchical progressive learning mechanism rooted in feature graph grain learning to mitigate potential conflict between memorization and forgetting in graph memory learning. This mechanism allows for a comprehensive and multi-level perception of local details within evolving graphs. In addition, to tackle the issue of unreliable structures in newly added incremental information, the paper introduces an information self-assessment ownership mechanism. This mechanism not only facilitates the propagation of incremental information within the model but also effectively preserves the integrity of past experiences. We design five types of graph memory learning tasks: regular, memory, unlearning, data-incremental, and class-incremental to evaluate BGML. Its excellent performance is confirmed through extensive experiments on multiple node classification datasets.
@@ I Introduction
Nowadays, a multitude of pivotal datasets in practical applications are expressed as graphs [1], encompassing diverse domains such as social networks [2, 3], traffic networks [4, 5], biological networks [6, 7], recommendation systems [8]. Graph learning models attract increasing attention and demonstrate robust data mining and representation learning capacities. However, graph learning faces the challenge of promptly assimilating new information while discarding obsolete knowledge that has been invalidated. This is attributed to the fact that real-world data often exhibits dynamic flow characteristics, wherein a continuous influx and withdrawal of data occur. As depicted in Fig. 1 (A), the flow of data across the graph frequently induces alterations in the overarching topology, thereby exacerbating the difficulties encountered by graph models. Consequently, the objective of this study is to explore and empower graph learning models with the capability to “memory” based on the dynamic graph. This paradigm is named as graph memory learning (GML).
@@
Existing methods in Graph Lifelong Learning (GLL) [9, 10] and Graph Unlearning (GUL) [11, 12] explore memory learning in graph models from a limited perspective. GLL mainly addresses the challenge of continuous data increments, aiming to adapt to new knowledge while mitigating the forgetting of past experiences [13, 14]. For instance, PI-GNN proposed by Zhang et al. [15] handles data increments through parameter isolation and expansion, while Su et al. [16] tackle catastrophic forgetting through structural alignment. However, GLL faces significant limitations when dealing with dynamic changes in graph data streams within the GML problem. Specifically, GLL fails to effectively adapt to significant changes in the local graph structure when data needs to be retracted. On the other hand, GUL emphasizes the right to forget data, aiming to forget specified retracted data at the lowest possible cost while minimizing harm to existing knowledge [17]. For example, the certified unlearning method proposed by Chien et al. [18] regularizes data retraction using convex function properties. Methods like IDEA [19] and ScaleGUN [20] have made improvements in flexibility and scalability. However, these methods primarily focus on static graph data and do not address dynamic graph learning scenarios within the GML problem, nor can they adapt to changes in the new knowledge environment based on forgetting. More deeply, GML is not simply a combination of GLL and GUL. It requires not only solving the challenges faced by them, but also a comprehensive and meticulous consideration of memory management in the context of continuous GML data flow. This is because the inherent conflict between memory retention and forgetting forms a natural barrier, preventing graph learning models from accurately perceiving changes in graph structure, thus leading to a decline in their expressive power. This conflict represents the most prominent contradiction in realizing the GML paradigm.
@@
In light of the above observations, this paper proposes a Brain-inspired Graph Memory Learning algorithm (BGML) based on brain network dynamics and function-structure coupling. It aims to solve the problem of GML for dynamically evolving graph data. Specifically, this paper proposes a multi-granular feature graph learning scheme with hierarchical progressive learning to clarify changes in local graph knowledge. Key components comprise a multi-granular graph partition (MGP) and a multi-granular hierarchical progressive learning mechanism (MGHPL) based on progress-aware module. Notably, progress-aware module evaluates and provides real-time feedback on the cognitive capacity of each sub-model, dynamically adjusting the integrated framework. Furthermore, an information self-assessment ownership mechanism (ISAO) is proposed to tackle the issue of unreliable new knowledge structures. Essentially, this paper explores three questions: (1) What GML approach can empower graph models to selectively remember important information and forget irrelevant information? (2) How to alleviate the conflict between remembering and forgetting in the graph memory model? (3) How to assign a credible graph structure to knowledge during the process of remembering and forgetting? The main contributions are summarized below:
@@
  • Extensive experiments on five tasks targeting different aspects of GML on the node classification task on nine datasets demonstrate that BGML outperforms all baselines significantly.
@@ II-A Graph Neural Networks
Graph neural networks (GNNs) have become a prominent method for processing unstructured data [24, 25, 26]. Specifically, Kpif et al. [27] introduced GCNs, utilizing first-order truncated Chebyshev polynomials to approximate graph convolutions. Additionally, Veličković et al. [28] introduced Graph Attention Networks (GATs), incorporating an attention mechanism on the graph, resulting in significant performance improvements. Furthermore, Hamilton et al. [29] were the first to identify the latent nodes within a graph, leading to the proposal of the inductive spatial graph neural network known as GraphSAGE. As research progresses, scholars have observed that the modeling of real-life scenes using graphs typically undergoes dynamic evolution, extending beyond static data [30, 31]. Based on distinct data forms (discrete and continuous), current research can be categorized into two groups [32]: discrete dynamic graph neural networks (DTDG), exemplified by DyTed [33] and SILD [34], and continuous dynamic graph neural networks (CTDG), exemplified by Ada-DyGNN [35] and SimpleDyG [36].
@@ II-B Lifelong Learning and Graph Lifelong Learning
Recently, these strategies have been progressively extended into the domain of graph learning [9]. Regarding architectural design, Wang et al. [13] devised the FGNs method to translate the graph structure into a format compatible with conventional data learning structures. To enhance the encoding of attribute information and the topological structure of target nodes, Zhang et al.’s [14] HPNs extract various levels of knowledge abstraction represented as atoms. TACO proposed by Han et al. [44] uses a topology-aware graph coarsening module to maintain the original graph topology information. Regarding the rehearsal scheme, Zhou et al. [45] store past knowledge in an experience buffer and conduct lifelong learning on graph data through experience replay. Su et al. [16] proposed SEA-ER by continuously using structural alignment to effectively combat catastrophic forgetting. In terms of regularization, Kou et al. [46] introduced decoupling-based continuous graph representation learning (DICGRL). Tan et al. [47] introduced GPIL with transferable meta-knowledge for initialization learning by introducing a pseudo-incremental approach. Furthermore, this field has given rise to several hybrid methods. ContinualGNN, proposed by Wang et al. [48], combines two approaches to alleviate catastrophic forgetting and preserve the existing learning mechanism. Feng et al. [49] proposed a triple replay architecture (OTGNet) to solve the problem of heterogeneous propagation. The SEM proposed by Zhang et al. [50] combines the advantages of replay and architecture and performs sparse sampling through Ricci curvature.
Existing GLL methods focus on addressing catastrophic forgetting, aiming to retain as much old experience as possible while adapting to new knowledge. However, with limited resources, continuous information accumulation leads to performance loss, and these methods struggle when specific knowledge must be forgotten. In contrast, BGML focuses on enabling the model to selectively remember important information while forgetting irrelevant information. This selective memory strategy allows the model to efficiently discard outdated information while conserving memory space, thereby enhancing its ability to continue learning.
@@ II-C Machine Unlearning and Graph Unlearning
To ensure data privacy and model safety, the machine must learn to forget [11, 12]. Cao et al. [51] were pioneers in introducing the concept of unlearning. Subsequently, researchers have proposed a series of new methods in two primary directions: exact unlearning and approximate unlearning [52, 53]. SISA (Sharded, Isolated, Sliced, and Aggregated), proposed by Bourtoule et al. [54], stands as one of the most representative works in this regard. The fundamental concept involves randomly dividing the training set into several disjoint shards and training the corresponding shard models independently. Upon receiving the forget request, the model only needs to retrain the respective shard model. Chatterjee et al. [55] discussed the possibility of unifying continual learning with machine unlearning.
Recently, Chen et al. [17] extended the concepts of SISA to the graph domain, presenting GraphEraser as the inaugural graph unlearning model. Almost at the same time, Chien et al. [18] used the convex function characteristics for regularization and proposed certified graph unlearning. Building on this, Dong et al. [19] (IDEA) and Yi et al. [20] (ScaleGUN) each focus on improving flexibility and scalability, respectively. Wu et al. [56] constructed the edge unlearning framework CEU. Differently, GNNDelete proposed by Cheng et al. [57] uses a model-independent hierarchical operator to optimize the deletion edge consistency and neighborhood influence for unlearning. Li et al. [58] introduced MEGU to chieve a balance between forgetting performance and framework generalization.
Current research on GUL remains primarily focused on static graphs, with limited exploration of how to adapt to environmental evolution during “unlearning”. BGML aims to apply this mechanism to dynamically evolving graphs, boldly addressing the memory issue in GML from the perspective of continuous increments and parallel forgetting.
@@ III Problem Formulation
The main target of GML is the dynamic evolution graph based on discrete snapshots, i.e., $\mathcal{G}=(\mathcal{G}_{0},\mathcal{G}_{1},\mathcal{G}_{2},...,\mathcal{G}_{t})$. The graph evolution process of each snapshot mainly includes two aspects. First, new data entering the graph network can be represented as $\mathcal{G}_{t}^{{}^{\prime}}=\mathcal{G}_{t-1}+\bigtriangleup\mathcal{G}_{t}^{{}^{\prime}}=\mathcal{G}_{t-1}\cup\left\{v_{\mathrm{new}},e_{\mathrm{new}}\mid v_{\mathrm{new}}\in\mathcal{V}_{\mathrm{add}}\right\}$, where $\mathcal{N}(v_{\mathrm{new}})$ represents the neighbors of $v_{\mathrm{new}}$ and $\mathcal{V}_{\mathrm{add}}$ represents the set of new nodes. Second, when the graph network receives forgetting requests (FR), the clearing operation can be expressed as $\mathcal{G}_{t}^{{}^{\prime\prime}}=\mathcal{G}_{t-1}+\bigtriangleup\mathcal{G}_{t}^{{}^{\prime\prime}}=\mathcal{G}_{t-1}\setminus\left\{v_{\mathrm{forget}},e_{\mathrm{forget}}\mid v_{\mathrm{forget}}\in\mathcal{V}_{\mathrm{delete}}\right\}$, where $\mathcal{V}_{\mathrm{delete}}$ is used to represent the set of nodes that need to be forgotten. In general, the change of the graph at time $t$ can be expressed as $\mathcal{G}_{t}=\mathcal{G}_{t-1}+\bigtriangleup\mathcal{G}_{t}$, where $\bigtriangleup\mathcal{G}_{t}=\bigtriangleup\mathcal{G}_{t}^{{}^{\prime}}+\bigtriangleup\mathcal{G}_{t}^{{}^{\prime\prime}}$. In addition, task context changes over time throughout the process of graph memory learning. The task set can be represented by $\mathcal{T}=(\mathrm{T}_{0},\mathrm{T}_{1},\mathrm{T}_{2},...,\mathrm{T}_{t})$. The corresponding time is $t=(t_{0},t_{1},t_{2},...,t_{t})$. Notely, given the complexity of GML, this paper specifically focuses on the problem using nodes as an example to maintain clarity and depth. The scenario where only edges change during graph evolution has not been discussed in this work.
@@ IV-B MGP: Multi-granular Graph Partition
Inspired by brain network principles such as the Sensory Pathways Theory and the Cortical Localization Theory (Remark 1), this paper proposes a multi-granularity graph partitioning approach. It mimics brain cognitive patterns to modularize and hierarchically decouple the graph knowledge.
Remark 1.
The Cortical Localization Theory posits that distinct regions of the cerebral cortex are specialized for processing different sensory, motor, and cognitive functions. The principles of Functional Localization and Regional Specialization suggest that specific areas of the cortex receive signals from particular sensory pathways. The Sensory Pathways Theory posits that sensory information is transmitted from external stimuli to the corresponding regions of the cerebral cortex for processing via specific neural pathways, such as the visual and auditory pathways. Furthermore, neuroimaging techniques (e.g. fMRI) have further confirmed that cortical division is precise and multi-level.
MGP Method. In conclusion, the cerebral cortex’s strategy of multi-region collaboration and hierarchical progression fosters the agent’s profound comprehension of knowledge. This necessitates preprocessing before knowledge input to facilitate problem division and resolution. Inspired by this, the paper devised an MGP algorithm, As shown in the MGP module in Fig. 2. This algorithm aims to partition the static regular graph $\mathcal{G}_{0}$ at time $t_{0}$ into multiple granules, establishing a hierarchy of knowledge details. Subsequent graph evolution details will be simultaneously incorporated into the granule. This lays the groundwork for the subsequent in-depth understanding and hierarchical extraction of precise information in the dynamic evolution graph. Specifically, the MGP algorithm is built upon the subgraph partitioning methods BLPA and BEKM proposed by Chen et al. [17]. For $\mathcal{G}_{0}$, the two graph partitions of coarse and fine granularity can be expressed by Eq.1:
@@
where $i=1,2,\dots,k$. $\mathcal{S}_{i}$ represents a first-level graph shard with coarse granularity. $\mathcal{P}_{j}^{i}$, on the other hand, denotes a second-level graph shard within $\mathcal{S}_{i}$, offering finer granularity, where $j=1,2,\dots,l$. $\mathrm{BLPA}(\cdot)$ and $\mathrm{BEKM}(\cdot)$ in Eq.1 are abstract representations of iterative algorithms (i.e. Eq.2 and Eq.3). Before iteration, they need to be provided with random initialized partition shards $\left\{\mathcal{S}_{1},\mathcal{S}_{2},\dots,\mathcal{S}_{k}\right\}^{0}$ and centroids $\left\{\mathbf{m}_{1},\mathbf{m}_{2},\dots,\mathbf{m}_{k}\right\}^{0}$, respectively.
For $\mathrm{BLPA}(\cdot)$, the core iteration steps in which $\forall v\in\mathcal{S}_{i}$ participates can be described as:
@@
$\mathrm{BEKM}(\cdot)$ performs clustering by distance between node embedding and centroids. For $\forall v$, the specific iteration steps can be expressed by Eq.3: